# FLASH: Latent-Aware Semi-Autoregressive Speculative Decoding for Multimodal Tasks

### v2

@@ 3.2 Visual token compression
Differ from LLM speculative decoding methods, our input consists of both visual and textual tokens. Compared to textual tokens, visual tokens are often numerous, leading to a significant increase in inference time Zhang et al. (2025). To balance the acceleration of draft generation and the potential degradation on acceptance rate, we propose an effective strategy for compressing the visual tokens in the draft model, allowing for faster inference without significantly compromising prediction quality. Since some visual tokens contain redundant information and do not significantly contribute to the final prediction, our goal is to compress the $N$ visual tokens into a smaller set of size $C$. Specifically, the $N$-sized visual feature $F_{V}$, corresponding to the $N$ visual tokens, is fed into a visual compressor to produce a compressed feature $\hat{F}_{V}$ of size $C$. Inspired by the attention mechanism, we introduce a learnable query set $\mathcal{C}$ of size $C$, where each query is designed to extract information from the feature $F_{V}$. By learning a compressed feature representation of size $C$, the model is able to retain the most salient semantics from the original visual feature. Specifically, the compressed feature $\hat{F}_{V}$ is computed as:
@@ 3.3 Semi-autoregressive inference
As Figure 2 shows, the semi-autoregressive head produces the hidden state feature $F^{\prime}_{M+1:M+K}$, which corresponds to the next $K$ tokens. These features are then projected into probability distributions over the vocabulary, denoted as $P^{\prime}_{1:K}$, using the frozen LM head identical to that of the target model. Finally, the candidate tokens $T^{\prime}_{1:K}$ are sampled from these distributions.
During training, our loss function comprises two components. The first is a regression loss, where we employ the Smooth L1 loss to quantify the difference between the features predicted by the draft model $F^{\prime}_{M+1:M+K}$ and those from the target model $F_{M+1:M+K}$, which serve as the ground truth:
$L_{1}=\sum_{i=M+1}^{M+K}\ell_{\text{reg}}(F^{\prime}_{i},F_{i}),$ (4)
where $\ell_{\text{reg}}(\cdot,\cdot)$ denotes the Smooth L1 loss. The second component is a classification loss, where we employ the cross-entropy loss to encourage the draft model to generate tokens that align with those produced by the target model:
$L_{2}=\sum_{i=1}^{K}\ell_{\text{cls}}(P^{\prime}_{i},P_{i}),$ (5)
where $\ell_{\text{cls}}(\cdot,\cdot)$ represents the cross-entropy loss, which is used to compare the probability distribution from the draft model and the target model. Therefore, the total loss is defined as:
$L=L_{1}+\alpha\cdot L_{2},$ (6)
where $\alpha$ is a hyper-parameter used to balance the regression and classification losses, ensuring they are of comparable magnitudes. Specifically, $\alpha$ is set to $0.1$. After computing the loss at each step, the next token predicted by the target model is appended to the input sequence. The extended sequence is then used to predict the subsequent $K$ tokens. This iterative procedure continues until the entire output sequence is generated.
@@
Algorithm 1 Draft Generation Process
1:Input: Visual embeddings $V$, Textual embeddings $E$, Number of tokens $K$, Semi-autoregressive head $S$, Compression query $\mathcal{C}$.
2:Output: Generated draft tokens ${T}^{\prime}_{1},...,{T}^{\prime}_{K}$.
3:Step 1: Transformer Processing
4:$F_{V},F_{T}\leftarrow\text{Transformer\_layers}(\text{concat}(V,E))$ // Second-to-top layer visual and textual features
5:Step 2: Visual token Compression
6:$\hat{F}_{V}\leftarrow\text{softmax}\left({\mathcal{C}\cdot F_{V}^{T}}\right)\cdot F_{V}$ // Compressed features
7:Step 3: Semi-autoregressive Generation
8:$\hat{F}_{T}\leftarrow\text{FC}(\text{concat}(F_{T},E_{N+2:M+1}))$ // Textual input
9:${F}^{\prime}\leftarrow\text{concat}(\hat{F}_{V},\hat{F}_{T})$ // Concatenate visual and textual input
10:${F}^{\prime}_{M+1},...,{F}^{\prime}_{M+K}\leftarrow S(F^{\prime})$ // Semi-autoregressive inference
11:${P}^{\prime}_{1},...,{P}^{\prime}_{K}\leftarrow\text{LM-head}(F^{\prime}_{M+1}),...,\text{LM-head}(F^{\prime}_{M+K})$ // Probability distribution of ${T}^{\prime}_{i}$
12:return ${T}^{\prime}_{1},...,{T}^{\prime}_{K}\leftarrow\text{argmax}({P}^{\prime}_{1}),...,\text{argmax}({P}^{\prime}_{K})$
@@ 4.1 Implementation settings
To evaluate the effectiveness of FLASH, we conduct experiments on LLaVA-1.5 and QwenVL-2.5 models. We train and evaluate on video captioning and visual instruction tuning task, using Kinetics-400 Kay et al. (2017) datasets and LLaVA-instruct-150k Liu et al. (2023), respectively. Specifically, Kinetics-400 is a video dataset containing 400 action classes. Since it only provides category annotations, we use the corresponding target model to generate captions as pseudo ground truth. LLaVA-Instruct-150K consists of multimodal instruction-response pairs generated by GPT. For both tasks, we sample 10k instances from each dataset for training. The maximum sequence lengths of training data are set to 200 for video captioning and 2048 for visual instruction tuning.
The learning rate is set to $2\times 10^{-5}$ and the batch size is set to 4. To ensure a fair comparison, all inferences are performed on a single NVIDIA A6000 GPU. During the evaluation, the batch size is set to 1, following standard practice in prior LLM speculative decoding methods Li et al. (2024b); Chen et al. (2024b).
We introduce two metrics to evaluate the effectiveness of the speculative decoding: the average acceptance tokens $\mathcal{A}$ and the speed-up ratio $\mathcal{R}$. The average acceptance tokens $\mathcal{A}$ is defined as the average number of the accepted tokens during a single forward pass of the draft model. Since multiple candidate tokens are generated in parallel during semi-autoregressive inference, the average acceptance tokens $\mathcal{A}$ can be greater than 1, which distinguishes it from autoregressive speculative decoding methods. The speed-up ratio $\mathcal{R}$ is defined as the generation speed of the target model divided by the total time required for token generation and verification under speculative decoding. We adopt speed-up ratio $\mathcal{R}$ as the primary metric to assess the overall efficiency.
Table 1: Quantitative experiments of video captioning on LLaVA-1.5 and QwenVL-2.5. We report the results of speed-up ratio $\mathcal{R}$ and the average acceptance tokens $\mathcal{A}$ under temperature $\tau=0$ and $\tau=1$. Additionally, we include the average computational cost (FLOPs), for a single round of draft generation and validation. FLASH combines “VisComp” and “SemiAR”, achieving the highest speed-up ratio among all competing methods, while maintaining output consistency with the target model. Underline indicates the best result, while green denotes the second-best result.
@@
Text-only 0.59 1.42×1.42$\times$1.42 × Text-only 0.52 1.37×1.37$\times$1.37 × 70.4T
Multimodal 0.69 1.63×1.63$\times$1.63 × Multimodal 0.66 1.60×1.60$\times$1.60 × 79.5T
@@
FLASH 2.632.63{{2.63}}2.63 $\underline{1.83\times}$ FLASH 2.632.63{2.63}2.63 $\underline{1.81\times}$ 70.7T
$\tau$=0 Speculative Decoding $\tau$=1 Speculative Decoding
Text-only 0.70 2.33×2.$3$ Text-only 0.54 1.61×1.61$\times$1.61 × 56.3T
Multimodal 0.83 2.49×2.49$\times$2.49 × Multimodal 0.80 1.93×1.93$\times$1.93 × 65.5T
Ours Ours
VisComp 0.83 2.60×2.60$\times$2.60 × VisComp 0.79 1.99×1.$9$ 57.3T
@@
FLASH 3.21 $\underline{2.68\times}$ FLASH 2.98 $\underline{2.05\times}$ 56.8T
4.2 Results
We perform experiments on two typical multimodal tasks: video captioning, which involves a large number of visual tokens as input, and visual instruction tuning, which processes a single image accompanied by textual interactions. Next, we analyze the experimental results for each task individually.
4.2.1 Video captioning
In video captioning, the primary challenge lies in effectively dispose the visual content while ensuring that the draft model can generate accurate captions. Among the competing models, “Text-only” refers to a speculative decoding method which only considers the textual input, following the approach proposed by  Gagrani et al. (2024). “Multimodal” refers to a variant of Eagle Li et al. (2024b), in which the original architecture is adapted to process multimodal input data. We observe that although text-only speculative decoding offers some acceleration, its speed-up ratio is inferior to that of multimodal speculative decoding. This is primarily because relying solely on textual context often causes the draft model to generate outputs that diverge from the actual visual content, thereby reducing the overall quality of the generated drafts.
In the LLaVA model, where each image initially consumes 576 tokens, our visual token compression method reduces this to 64 tokens. Comparing “Multimodal” and “VisComp”, we observe that incorporating vision token compression significantly reduces the computational load (FLOPs) and improves the speed-up ratio $\mathcal{R}$ by reducing the number of input tokens, which causing only a minimal decrease in the average acceptance tokens $\mathcal{A}$. Compared to speculative decoding in the autoregressive setting (“Multimodal”), the semi-autoregressive approach (“SemiAR”) demonstrates a significant advantage in terms of average accepted tokens, as it can return multiple draft candidates in a single forward pass. This leads to an overall speed-up ratio improvement of approximately $9.3\%$. By combining visual token compression with semi-autoregressive inference, FLASH achieves the highest speed-up ratio at about 1.8×1.8$\times$1.8 × compared to the target model.
In the QwenVL model, the number of visual tokens varies with the input image size. To standardize, we resize images to generate 324 tokens per image. Additionally, our visual token compression reduces the token count to 36. Most trends in the QwenVL model align with those observed in LLaVA. Compared to “Multimodal”, “VisComp” and “SemiAR” yield additional speed-up gains of approximately $2.6\%$ and $6.3\%$, respectively. Furthermore, FLASH achieves speed-up ratios of 2.68×2.68$\times$2.68 × and 2.05×2.05$\times$2.05 × over target model under $\tau=0$ and $\tau=1$, respectively. These results demonstrate the effectiveness of our method on both LLaVA and QwenVL models.
Table 2: Quantitative experiments of visual instruction tuning on LLaVA-1.5 and QwenVL-2.5. Similar to video captioning task, we report speed-up ratio $\mathcal{R}$, average acceptance tokens $\mathcal{A}$, and average computational cost (FLOPs). FLASH combines “VisComp” and “SemiAR”, achieving the highest speed-up ratio among all competing methods, while maintaining output consistency with the target model. Underline indicates the best result, while green denotes the second-best result.
$\tau$=0 Speculative Decoding $\tau$=1 Speculative Decoding
Text-only 0.68 1.60×1.60$\times$1.60 × Text-only 0.66 1.52×1.52$\times$1.52 × 11.6T
Multimodal 0.76 2.21×2.21$\times$2.21 × Multimodal 0.76 2.19×2.19$\times$2.19 × 12.7T
Ours Ours
VisComp 0.76 2.23×2.23$\times$2.23 × VisComp 0.76 2.21×2.21$\times$2.21 × 11.8T
@@ 4.2.1 Video captioning
FLASH 2.77 $\underline{2.55\times}$ FLASH $\underline{2.59}$ $\underline{2.29\times}$ 11.5T
$\tau$=0 Speculative Decoding $\tau$=1 Speculative Decoding
Text-only 0.52 1.44×1.$4$ Text-only 0.50 1.36×1.36$\times$1.36 × 11.4T
Multimodal 0.65 1.63×1.63$\times$1.63 × Multimodal 0.63 1.59×1.59$\times$1.59 × 12.6T
Ours Ours
VisComp 0.63 1.69×1.69$\times$1.69 × VisComp 0.60 1.65×1.65$\times$1.65 × 11.6T

### v3

@@ 3.2 Visual token compression
Unlike LLM speculative decoding methods, our input consists of both visual and textual tokens. Compared to textual tokens, visual tokens are often numerous, leading to a significant increase in inference time (Zhang et al., 2025). To balance the acceleration of draft generation and the potential degradation on acceptance rate, we propose an effective strategy for compressing the visual tokens in the draft model, allowing for faster inference without significantly compromising prediction quality. Since some visual tokens contain redundant information and do not significantly contribute to the final prediction, our goal is to compress the $N$ visual tokens into a smaller set of size $C$. Specifically, the $N$-sized visual feature $F_{V}$, corresponding to the $N$ visual tokens, is fed into a visual compressor to produce a compressed feature $\hat{F}_{V}$ of size $C$. Inspired by the attention mechanism, we introduce a learnable query set $\mathcal{C}$ of size $C$, where each query is designed to extract information from the feature $F_{V}$. By learning a compressed feature representation of size $C$, the model is able to retain the most salient semantics from the original visual feature. Specifically, the compressed feature $\hat{F}_{V}$ is computed as:
@@ 3.3 Semi-autoregressive inference
To enable parallel draft decoding, we input $K$ placeholder tokens to denote the future positions without exposing future target tokens. During verification, we apply the standard token-wise acceptance–rejection rule using the draft probabilities produced for each position and the target probabilities computed in parallel by the target model (Leviathan et al., 2023). As Figure 2 shows, the semi-autoregressive head outputs hidden states $F^{\prime}_{M+1:M+K}$, which are projected by the frozen LM head into probability distributions $P^{\prime}_{1:K}$, from which candidate tokens $T^{\prime}_{1:K}$ are sampled.
Our loss function comprises two components. The first is a regression loss, where we employ the Smooth L1 loss to quantify the difference between the features predicted by the draft model $F^{\prime}_{M+1:M+K}$ and those from the target model $F_{M+1:M+K}$. The second component is a classification loss, where we employ the cross-entropy loss to encourage the draft model to generate tokens that align with those produced by the target model. The total loss is defined as:
$L=\sum_{i=M+1}^{M+K}\ell_{\text{reg}}(F^{\prime}_{i},F_{i})+\alpha\cdot\sum_{i=1}^{K}\ell_{\text{cls}}(P^{\prime}_{i},P_{i}),$ (4)
where $\ell_{\text{reg}}(\cdot,\cdot)$ and $\ell_{\text{cls}}(\cdot,\cdot)$ denotes the Smooth L1 loss and the cross-entropy loss, respectively. $\alpha$ is a hyper-parameter used to balance the regression and classification losses, ensuring they are of comparable magnitudes. Specifically, $\alpha$ is set to $0.1$. After computing the loss at each step, the next token predicted by the target model is appended to the input sequence. The extended sequence is then used to predict the subsequent $K$ tokens. This iterative procedure continues until the entire output sequence is generated.
@@ 4.1 Implementation settings
We conduct experiments using LLaVA-1.5 and QwenVL-2.5 as target models. We train and evaluate on video captioning (VC) and visual instruction tuning (VIT), using Kinetics-400 (Kay et al., 2017) and LLaVA-instruct-150k (Liu et al., 2023), respectively. We use Kinetics-400 mainly as an efficiency stress test with long visual contexts. Since our goal is to accelerate target LMM without changing its output distribution, we use target-generated captions to train the draft model to match the target behavior under the same prompts. For both tasks, we sample 10k instances from each dataset for training. The maximum sequence lengths of training data are set to 200 for VC and 2048 for VIT.
The learning rate is set to $2\times 10^{-5}$ and the batch size is set to 8 in our main results. All inferences are performed on a single NVIDIA A6000 GPU with batch size set to 1, following standard practice in prior LLM speculative decoding methods (Li et al., 2024c; Chen et al., 2024c).
We introduce two metrics to evaluate the effectiveness of the speculative decoding: the average acceptance tokens $\mathcal{A}$ and the speed-up ratio $\mathcal{R}$. The average acceptance tokens $\mathcal{A}$ is defined as the average number of the accepted tokens during a single forward pass of the draft model. Since multiple candidate tokens are generated in parallel during semi-autoregressive inference, the average acceptance tokens $\mathcal{A}$ can be greater than 1, distinguishing it from autoregressive speculative decoding methods. The speed-up ratio $\mathcal{R}$ is adopted as the primary metric to assess the overall efficiency, defined as the generation time of the target model divided by the total time required for speculative decoding including the draft token generation and verification.
Table 1: Quantitative experiments of video captioning on LLaVA-1.5 and QwenVL-2.5. We report the results of speed-up ratio $\mathcal{R}$ and the average acceptance tokens $\mathcal{A}$ under temperature $\tau=0$ and $\tau=1$. Additionally, we include the average computational cost (FLOPs), for a single round of draft generation and validation. SpecFLASH combines “VisComp” and “SemiAR”, achieving the best speed-up ratio among all competing methods, while maintaining output consistency with the target model. Underline indicates the best result, while green denotes the second-best result.
@@
Text-only (Gagrani et al., 2024) 0.59 $1.42\times$ Text-only (Gagrani et al., 2024) 0.52 $1.37\times$ 70.4T
Eagle-MM (Li et al., 2024c) 0.69 $1.63\times$ Eagle-MM (Li et al., 2024c) 0.66 $1.60\times$ 79.5T
Medusa-MM (Cai et al., 2024) 2.31 $1.69\times$ Medusa-MM (Cai et al., 2024) 2.27 $1.63\times$ 80.2T
SpecVLM (Huang et al., 2025) 0.62 $1.46\times$ SpecVLM (Huang et al., 2025) 0.59 $1.40\times$ 81.0T
Dream (Hu et al., 2025) 0.70 $1.66\times$ Dream (Hu et al., 2025) 0.68 $1.63\times$ 79.9T
@@
VisComp 0.68 $1.79\times$} VisComp 0.66 $1.70\times$ 71.2T
SemiAR $\underline{2.65}$ $1.77\times$ SemiAR $\underline{2.64}$ $1.76\times$} 74.5T
SpecFLASH 2.63{{2.63}} $\underline{1.83\times}$ SpecFLASH 2.63{2.63} $\underline{1.81\times}$ 70.7T
$\tau$=0 Speculative Decoding $\tau$=1 Speculative Decoding
Text-only (Gagrani et al., 2024) 0.70 $2.33\times$ Text-only (Gagrani et al., 2024) 0.54 $1.61\times$ 56.3T
Eagle-MM (Li et al., 2024c) 0.83 $2.49\times$ Eagle-MM (Li et al., 2024c) 0.80 $1.93\times$ 65.5T
Medusa-MM (Cai et al., 2024) 2.76 $2.41\times$ Medusa-MM (Cai et al., 2024) 2.65 $\underline{2.35\times}$ 67.3T
SpecVLM (Huang et al., 2025) 0.64 $1.62\times$ SpecVLM (Huang et al., 2025) 0.63 $1.58\times$ 79.3T
Dream (Hu et al., 2025) 0.83 $2.51\times$ Dream (Hu et al., 2025) 0.82 $2.00\times$ 65.7T
Ours Ours
VisComp 0.83 $2.60\times$ VisComp 0.79 $1.99\times$ 57.3T
SemiAR $\underline{3.28}$ $2.63\times$} SemiAR $\underline{3.09}$ $2.00\times$} 61.6T
SpecFLASH 3.21 $\underline{2.68\times}$ SpecFLASH 2.98 $2.05\times$ 56.8T
4.2 Results
We perform experiments on two typical multimodal tasks: video captioning, which involves a large number of visual tokens as input, and visual instruction tuning, which processes a single image accompanied by textual interactions.
4.2.1 Video captioning
In video captioning (VC), the primary challenge lies in effectively processing the visual content while ensuring that the draft model can generate accurate captions. Among the competing models, “Text-only” refers to a speculative decoding method which does not consume visual embeddings and does not generate multimodal-conditioned draft predictions, following the approach proposed by  (Gagrani et al., 2024). “Eagle-MM” and “Medusa-MM” refer to variants of Eagle (Li et al., 2024c) and Medusa (Cai et al., 2024), to process image inputs by incorporating a CLIP-style visual encoder and projector, which align visual and textual embeddings into a shared space. These aligned embeddings are concatenated to form the input embeddings and are combined with the second-to-top layer feature $F$, following the same design as used in the textual components. We observe that although text-only speculative decoding offers some acceleration, its speed-up ratio is inferior to that of multimodal speculative decoding. This is primarily because relying solely on textual context often causes the draft model to generate outputs that diverge from the actual visual content, thereby reducing the overall quality of the generated drafts.
In the LLaVA model, where each image initially consumes 576 tokens, our visual token compression method reduces this to 64 tokens. Comparing previous speculative decoding method “Eagle-MM” and “Medusa-MM” with “VisComp”, we observe that incorporating visual token compression significantly reduces the computational load (FLOPs) and improves the speed-up ratio $\mathcal{R}$ by reducing the number of input tokens, which causes only a minimal decrease in the average acceptance tokens $\mathcal{A}$. Compared to speculative decoding in the autoregressive setting (“Eagle-MM”), the semi-autoregressive approach (“SemiAR”) demonstrates a significant advantage in terms of average accepted tokens, as it can return multiple draft candidates in a single forward pass. This leads to an overall speed-up ratio improvement of approximately $9.3\%$. By combining visual token compression with semi-autoregressive inference, SpecFLASH achieves the highest speed-up ratio at about $1.8\times$ compared to the target model.
Table 2: Quantitative experiments of visual instruction tuning on LLaVA-1.5 and QwenVL-2.5. Similar to VC task, we report speed-up ratio $\mathcal{R}$, average acceptance tokens $\mathcal{A}$, and average computational cost (FLOPs). SpecFLASH combines “VisComp” and “SemiAR”, achieving the highest speed-up ratio among all competing methods, while maintaining output consistency with the target model. Underline indicates the best result, while green denotes the second-best result.
$\tau$=0 Speculative Decoding $\tau$=1 Speculative Decoding
Text-only (Gagrani et al., 2024) 0.68 $1.60\times$ Text-only (Gagrani et al., 2024) 0.66 $52\times 11.$6T
Eagle-MM (Li et al., 2024c) 0.76 $2.21\times$ Eagle-MM (Li et al., 2024c) 0.76 $2.19\times$ 12.7T
Medusa-MM (Cai et al., 2024) 2.55 $2.28\times$ Medusa-MM (Cai et al., 2024) 2.52 $2.24\times$ 13.0T
SpecVLM (Huang et al., 2025) 0.73 $1.77\times$ SpecVLM (Huang et al., 2025) 0.69 $1.72\times$ 13.2T
Dream (Hu et al., 2025) 0.72 $2.18\times$ Dream (Hu et al., 2025) 0.70 $2.14\times$ 12.9T
Ours Ours
VisComp 0.76 $2.23\times$ VisComp 0.76 $2.21\times$ 11.8T
SemiAR $\underline{2.86}$ $2.49\times$} SemiAR 2.58 $2.24\times$} 12.0T
SpecFLASH 2.77 $\underline{2.55\times}$ SpecFLASH $\underline{2.59}$ $\underline{2.29\times}$ 11.5T
$\tau$=0 Speculative Decoding $\tau$=1 Speculative Decoding
Text-only (Gagrani et al., 2024) 0.52 $1.44\times$ Text-only (Gagrani et al., 2024) 0.50 $36\times 11.$4T
Eagle-MM (Li et al., 2024c) 0.65 $1.63\times$ Eagle-MM (Li et al., 2024c) 0.63 $1.59\times$ 12.6T
Medusa-MM (Cai et al., 2024) 2.36 $1.68\times$ Medusa-MM (Cai et al., 2024) 2.29 $1.65\times$ 13.0T
SpecVLM (Huang et al., 2025) 0.65 $1.52\times$ SpecVLM (Huang et al., 2025) 0.62 $1.50\times$ 13.2T
Dream (Hu et al., 2025) 0.67 $1.66\times$ Dream (Hu et al., 2025) 0.66 $1.65\times$ 12.7T
Ours Ours
VisComp 0.63 $1.69\times$ VisComp 0.60 $65\times 11.$6T
SemiAR 2.46 $1.79\times$} SemiAR $\underline{2.38}$ $1.70\times$} 11.8T
SpecFLASH $\underline{2.46}$ $\underline{1.83\times}$ SpecFLASH 2.36 $\underline{1.71\times}$ 11.3T
In the QwenVL model, the number of visual tokens varies with the input image size. To standardize, we resize images to generate 324 tokens per image. Additionally, our visual token compression reduces the token count to 36. Most trends in the QwenVL model align with those observed in LLaVA. Compared to “Eagle-MM”, “VisComp” and “SemiAR” yield additional speed-up gains of approximately $2.6\%$ and $6.3\%$, respectively. Furthermore, SpecFLASH achieves speed-up ratios of $2.68\times$ and $2.05\times$ over target model under $\tau=0$ and $\tau=1$, respectively. These results demonstrate the effectiveness of our method on both LLaVA and QwenVL models.
4.2.2 Visual instruction tuning
In visual instruction tuning (VIT), each input consists of an image paired with a textual instruction, covering a broad range of tasks such as image captioning, visual question answering, object recognition, and visual reasoning. Using LLaVA-1.5 and QwenVL-2.5 as the target models, we conduct experiments on VIT task following the same experiment setups as in VC task.
As presented in Table 2, by comparing the “Text-only” with “Eagle-MM” speculative decoding, we find that relying solely on text input leads to a significant drop in the draft generation quality when fine-grained image understanding is required, particularly on LLaVA. Additionally, our proposed visual token compression (“VisComp”) and semi-autoregressive (“SemiAR”) improves the overall speed-up ratio by $2.3\%$ and $7.9\%$, respectively, comparing to “Eagle-MM”. Specifically, we further find that when temperature $\tau=0$, using “SemiAR” yields a $0.22\times$ speed-up, which is notably higher than the $0.08\times$ improvement observed at $\tau=1$. This improvement due to the model’s tendency to produce more deterministic outputs at lower temperatures, which aligns well with the parallel token generation in the semi-autoregressive head.
Notably, “SemiAR” yields greater improvements in VIT than in VC. Besides, we observe that the improvement brought by visual token compression is less pronounced in the visual instruction tuning task compared to VC task. This is likely because visual instruction tuning typically involves a single image, whereas VC tasks process multiple frames. By combining “VisComp” and “SemiAR”, SpecFLASH achieves speed-up ratios of $2.55\times$ and $1.83\times$ on LLaVA and QwenVL, respectively. Moreover, we observe that SpecFLASH efficiently processes visual information while incurring a computational load comparable to that of the “Text-only” model. It enables SpecFLASH to achieve a significantly higher speed-up ratio without sacrificing draft quality. These results further demonstrate the effectiveness of our method across both VC and VIT tasks.
4.2.3 Text-only Scenario
To further investigate the effectiveness of our method beyond multimodal tasks, we additionally evaluate SpecFLASH in text-only setting. In this case, the visual token compression module becomes inactive, and the draft model operates solely based on the textual prompt.
Table 3: Speed-up ratios under text-only prompts on LLaVA-650k subset.
Method Image-text pairs Text-only
Text-only - $1.63\times$
Eagle-$M$ $2.21\times$ $1.82\times$
SpecFLASH $2.49\times$ $1.88\times$
The results on LLaVA-650k text-only subset are summarized in Table 3. Several observations can be drawn. First, speculative decoding in a text-only setting achieves a $1.63\times$ speed-up, which is lower than that in the multimodal setting. The results indicate that visual input can provide structural cues (e.g., spatial relations) that tend to make generation more predictable. In comparison, text-only sequences generally display more diverse linguistic patterns. Second, SpecFLASH attains $1.88\times$ speed-up in the text-only case ($0.25\times$ over the text-only baseline), and $2.49\times$ in the multimodal setting ($0.28\times$ over Eagle-MM).
Overall, these results confirm two key insights: First, visual grounding indeed provides additional structural constraints that amplify the benefits of speculative decoding. Second, SpecFLASH’s semi-autoregressive design and draft architecture remain effective in pure-text scenarios, even though the model is primarily trained on image–text pairs. We hypothesize that explicit exposure to text-only data during training could further boost performance in such cases. These findings highlight the distinct structural properties of multimodal generation and reinforce our claim that semi-autoregressive decoding is particularly well-suited to vision–language models.
4.3 Ablations
To further clarify the contribution of the visual token compression module, we perform an ablation where SpecFLASH is evaluated with different visual compression methods. Results are shown in Table 4.
Table 4: Ablation study on different visual token compression methods.
Method VIT VC
w/o VisComp $2.49\times$ $1.77\times$
Average pooling $2.31\times$ $1.77\times$
SGL (Zhao et al., 2025) $1.68\times$ $1.80\times$
FastV (Chen et al., 2024a) $1.53\times$ $1.60\times$
Twig (Shao et al., 2025) $1.58\times$ $1.53\times$
SpecFLASH (ours) $2.55\times$} $1.83\times$}
We observe that our SpecFLASH achieves the best acceleration consistently across both tasks ($2.55\times$ and $1.83\times$). Non-parametric strategies such as average pooling and SGL (Zhao et al., 2025) yield limited speed-up ratio. It suggests that visual token compression module goes beyond simple redundancy-reduction heuristics, providing a more principled design to maximize efficiency. In contrast to previous visual token compression methods like FastV (Chen et al., 2024a) and Twig (Shao et al., 2025), which primarily aim to shorten the input sequence and preserve semantic information, our approach emphasizes inference efficiency. As a result, while their acceleration is limited, it reflects the fact that their design focus was not faster decoding but semantic preservation under compression.
To further explore the trade-off between the draft quality and its efficiency, we vary the scale of the draft model through adjusting the layer parameters. To be specific, we modify the number of parameters in the draft model, including the number of attention heads and the number of attention layers. The resulting speed-up ratios are reported in Table 5. These results show that balance between the scale of draft model and speed is important. For example, a draft model with 2 Transformer layers may have more accuracy in predicting the target distributions. However, at the same time, it introduces additional computational cost thus affect the overall speed-up ratios.
Table 5: Speed-up ratios under different draft model scales.
Task VIT VC
@@ 4.3 Ablations
5 Conclusions
In this work, we introduce SpecFLASH, a novel speculative decoding framework designed specifically for Large Multimodal Models (LMMs). SpecFLASH introduces two key components: a visual token compression module that reduces input redundancy, and a semi-autoregressive decoding strategy that enables parallel token generation. Together, these innovations accelerate multimodal inference while preserving the quality of the output. Experiments on video captioning and visual instruction tuning tasks demonstrate that SpecFLASH delivers substantial speed-ups over existing speculative decoding methods, enabling more efficient deployment of LMMs.
Appendix A Scale of Draft Model
Other than changing the scale of the draft model by adjusting the number of parameters, we conducted additional experiments by applying quantization to the draft model. The measured speed-up ratios are summarized in Table 6.
Table 6: Quantization on draft model, reporting the speed-up ratio as the primary metric.
Task VIT VC
SpecFLASH-fp16 $2.55\times$ $1.83\times$
SpecFLASH-bf16 $2.43\times$ $1.79\times$
SpecFLASH-int8 $2.44\times$ $1.74\times$
SpecFLASH-int4 $2.15\times$ $1.55\times$
These results show that quantization does not offer the expected further speeding up. We believe this is because our draft model is already lightweight. In such a small model, the efficiency gains from quantization are insufficient to offset the loss in prediction accuracy.
Appendix B Algorithm
Algorithm 1 illustrates the draft generation process. Code is provided in Supplementary Materials.
Algorithm 1 Draft Generation Process
1:Input: Visual embeddings $V$, textual embeddings $E$, number of tokens $K$, semi-autoregressive head $S$, compression query $\mathcal{C}$
2:Output: Draft tokens $T^{\prime}_{1},\dots,T^{\prime}_{K}$
3: // Transformer processing